# High-OAM Deep Ultraviolet Twisted Light Generation for RF-Photoinjector Applicationsthanks: This document is the results of the research project funded by the Russian Science Foundation (Project No. 23-62-10026) [RSF]

### v3

@@ 4.1 Experimental Setup
Figure 9: Schematic layout of the experimental setup. The system includes the RF photoinjector drive laser; periscope mirrors M1–M2 (beam extraction) and M3–M4 (beam transfer); spatial-filtering lens L1 and pinhole PH; optional lens L2 for fork-grating measurements; the tested phase element (fork grating, SPP, or axicon); a cylindrical lens for the OAM diagnostics; and a UV-sensitive CCD camera for beam imaging.
After spatial filtering, the beam illuminated the DOE under test (fork grating, SPP, or axicon). For the OAM diagnostics in the fork-grating and SPP experiments, a cylindrical lens with focal length $f_{\mathrm{cyl}}=100\,$mm was used to perform Laguerre–Gaussian to Hermite–Gaussian (LG-to-HG) mode conversion [Beijersbergen92, Kotlyar17]. The resulting intensity patterns were recorded using a UV-sensitive SDU-285R CCD camera.
@@ 4.2 Experiments with Fork Grating
The reflective fork grating was first tested using the UV beam prepared by the spatial-filtering stage. The generated twisted light are shown in Fig. 11. In these measurements, the central (zeroth-order) beam was blocked to avoid saturation of the CCD sensor. As expected for a grating with $m=2$, the $\pm 1$ and $\pm 2$ diffraction orders produced Laguerre–Gaussian modes with orbital angular momenta $\ell=\pm 2$ and $\ell=\pm 4$, respectively.
Power measurements using an Ophir 3A-FS detector confirmed that the majority of the transmitted energy was contained in the zeroth and first diffraction orders, while higher orders carried progressively smaller power fractions.
The corresponding Hermite–Gaussian modes obtained after cylindrical-lens conversion are presented in Fig. 11(b). The observed modal structures agree with the expected relation [Beijersbergen92, Kotlyar17]
@@
demonstrating that the grating reliably generates the targeted OAM states. Higher diffraction orders (n=3n=-3 and n=4n=-4), shown in Fig. 11, also exhibit well-defined modal patterns, confirming correct operation of the fabricated structure across multiple orders.
Figure 10: Twisted light generated by the reflective fork grating for diffraction orders $\pm 1$ and $\pm 2$. (a) Laguerre–Gaussian modes with the OAM $\ell=\pm 2$ and $\ell=\pm 4$. (b) Corresponding Hermite–Gaussian modes obtained after cylindrical-lens mode conversion.
Figure 11: Higher-order twisted light generated by the reflective fork grating. (a) Laguerre–Gaussian modes with the OAM $\ell=-2,\ -4,\ -6,\ -8$ observed in the n=1n=-1, 2-2, 3-3, and 4-4 diffraction orders. (b) Corresponding Hermite–Gaussian modes obtained after cylindrical-lens mode conversion.
A potential route for improving the performance of the diffractive element is to increase the spatial resolution of the grating lines. A finer grating would enable generation of higher-quality vortex modes but would also require tighter focusing of the incident beam onto the grating surface. Overall, the experimentally measured spatial profiles in Fig. 11 show good qualitative agreement with the simulations presented in Fig. 6, confirming correct operation of the fabricated fork grating.
@@ 4.3 Experiments with SPP
The next experiment was performed using the fabricated SPP, and the resulting beam profiles are shown in Fig. 13. The output twisted light (Fig. 13(a)) exhibits a well-defined annular structure with a pronounced central singularity. A portion of the beam energy is redistributed into several weak concentric rings surrounding the primary vortex mode, consistent with the expected behavior of high-charge SPP-generated beams.
Power measurements performed with the Ophir 3A-FS detector showed an intensity reduction of approximately $20\%$. Although this differs from the theoretical conversion efficiency of $\sim 90\%$, the obtained value is still reasonable given the surface-profile discretization, potential scattering losses, and alignment constraints in the UV. The measured twisted light structure and transverse size closely match the simulated pattern shown in Fig. 8.
The Hermite–Gaussian mode exhibits a slightly larger tilt compared with the simulation. This effect is attributed to a small intentional misalignment of the beam relative to the focal plane of the cylindrical lens, introduced to improve the separation and visibility of the experimentally observed mode stripes.
4.4 Experiments with Axicons
The final series of experiments was performed using the fabricated binary axicons. The resulting quasi–Bessel beams are shown in Fig. 13. For the axicon with the topological charge $m=3$ (Fig. 13(a)), three intensity lobes are clearly visible, consistent with the expected topological charge. At short propagation distances, these lobes partially overlap and form a characteristic Y-shaped pattern, which is typical for binary axicons due to the limited spatial separation between the initial diffraction maxima.
Figure 12: Twisted light generated by the SPP. (a) Laguerre–Gaussian vortex carrying the OAM $\ell=64$. (b) Corresponding Hermite–Gaussian mode of the order $N=65$ obtained after cylindrical-lens mode conversion.
Figure 13: Twisted light generated by the binary axicons. (a) Quasi–Bessel beam produced by the axicon with the topological charge $m=3$, exhibiting three intensity lobes. (b) Quasi–Bessel beam produced by the axicon with the topological charge $m=10$, exhibiting ten resolved lobes.
In contrast, the axicon with $m=10$ produces a well-resolved multi-lobe ring (Fig. 13(b)), in good agreement with the simulated patterns in Fig. 8. Both axicons exhibit weak spiral-like background features associated with the binary phase discretization.
@@ 4.4 Experiments with Axicons
Finally, we note that Bessel and quasi–Bessel beams cannot be converted into Hermite–Gaussian modes using a cylindrical-lens scheme, since their propagation-invariant structure does not undergo the conventional LG-to-HG transformation. For such beams, the orbital angular momentum projection simply coincides with the topological charge $m$. As shown in Ref. [Knyazev23], $m$ can be obtained simply by counting the number of intensity maxima in the ring, providing a straightforward and reliable method for identifying the OAM content of quasi–Bessel beams. In practice, quasi–Bessel beams can exhibit a finite OAM bandwidth due to slight contributions from nearby OAM states, but the average topological charge still corresponds to $m$, as discussed in Section 3.
@@ 5 Discussion and Conclusion
Figure 14: Radial profiles of the generated twisted ligh beams: (a)(c) – fork grating producing the OAM $\ell=2$, $\ell=4$ and $\ell=8$ modes; (d) – SPP generating a high-order LG beam with $\ell=64$; (e) – binary axicon producing a quasi–Bessel beam with the topological charge $m=10$.
We have demonstrated the generation of high-OAM ultraviolet twisted light using three types of diffractive optical elements: a fork grating, a spiral phase plate, and binary axicons. The fork grating, being the simplest and most cost-effective DOE, produced vortex modes of high quality (Fig. 14(a)–(c)) with well-formed single-ring structures and clear central minima. As discussed in Sec. 4.2, the first ring width reflects the conversion efficiency from the incident Gaussian beam: lower topological charges yield broader rings, while higher charges produce narrower rings that gradually widen with increasing OAM, consistent with LG modal scaling. The ring radius increase systematically with OAM, matching theoretical predictions (Fig. 6).
Measured energy distribution indicates a substantial fraction of power in zeroth and low diffraction orders, limiting overall conversion efficiency. Nevertheless, stable generation of $\ell=2$, $\ell=4$, and $\ell=8$ OAM states with good control of LG modal structure was achieved.
The grating period could be reduced to improve efficiency, but in our setup this is constrained by the resolution of the laser-engraving system. Fork gratings remain scalable to higher charges ($m\sim 8$–10), enabling access to $\ell\sim 10$–20 in low diffraction orders, making them suitable for semiconductor photocathodes (e.g., Cs2Te, GaAs) with high quantum efficiency [Xiang15].
The SPP showed the best overall performance. Its radial profile (Fig. 14(d)) exhibits a smooth annular intensity distribution with a deep central minimum and weak secondary rings, confirming excellent modal purity and high conversion efficiency ($\approx 80\%$). This enabled generation of an $\ell=64$ beam with well-defined transverse structure suitable for OAM transfer to relativistic electrons and RF-photoinjector diagnostics. Since the divergence of an LG mode scales as $\sqrt{\ell}$ [Karlovets21], beamlines intended for high-OAM injection should take into account the correspondingly larger Twiss $\beta$-function [Wiedemann15, Chao20, Lee21] and transverse emittance. The high modal purity of the SPP beam also facilitates OAM diagnostics using LG–HG conversion with astigmatic optics [Schattschneider12].
Binary axicons produced quasi–Bessel beams with $\approx 70\%$ efficiency and robust spatial structure. Despite the multi-lobe transverse pattern typical for binary axicons (Fig. 13(a)), clean radial intensity distributions were observed (Fig. 14(e)), in agreement with topological charge $m=10$ and numerical simulations in Sec. 3.
For RF-photoinjector implementation with metallic photocathodes, SPPs and axicons are most attractive due to higher efficiency, better modal purity, and superior power handling. All beams exhibit transverse feature sizes of one to several hundred micrometers, enabling high–charge-density, high-brightness electron beams critical for synchrotrons and free-electron lasers. Diffraction on a triangular aperture provides a promising method to determine OAM for quasi–Bessel beams [Maksimov25], as LG–HG conversion is inapplicable. Table 1 summarizes our key results and recommendations.
Table 1: Comparison of parameters for a fork grating, spiral phase plate, and binary axicon.
@@
Aknowledgement
We deeply acknowledge A.A. Bogdanov for the help that made the fabrication of the spiral phase plate possible. We sincerely thank K.V. Cherepanov for invaluable assistance and advice. We are also grateful to S.S. Baturin for helpful discussions and constructive feedback. We additionally thank N.E. Sheremet for assistance with developing the numerical code used to generate the fork-grating phase masks. The axicons were simulated and fabricated at the shared research facility of Siberian Center for Synchrotron and Terahertz Radiation at the basis of the Novosibirsk Free Electron Laser, Budker Institute of Nuclear Physics SB RAS.
This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [RSF].

### v4

@@ 4.1 Experimental setup
Figure 12: Schematic layout of the experimental setup and representative beam-formation configurations. (a) General view of the optical experiment, including the drive laser beamline, beam-conditioning optics, and the detection system. (b) Diffraction of the laser beam by a reflective fork grating, yielding Laguerre–Gaussian modes in several diffraction orders with topological charges proportional to the grating topological charge $m=2$. (c) Generation of twisted light carrying orbital angular momentum $\ell=64\hbar$ using a spiral phase plate (SPP). (d) Formation of a coherent superposition of quasi–Bessel twisted light modes produced by a binary axicon.
@@
After spatial filtering, the beam illuminated the DOE under test (fork grating, SPP, or axicon). For the OAM diagnostics in the fork-grating and SPP experiments, a cylindrical lens with a focal length of $100$ mm was used to perform Laguerre–Gaussian to Hermite–Gaussian (LG-to-HG) mode conversion [6, 29]. The resulting intensity distributions were recorded with a UV-sensitive SDU-285R CCD camera.
@@ 4.2 Experiments with fork grating
The reflective fork grating was first tested using the UV beam prepared by the spatial-filtering stage. The generated twisted light beams are shown in Figure 13. In these measurements, the central (zeroth-order) beam was blocked to avoid saturation of the CCD sensor. As expected for a grating with $m=2$, the $\pm 1$ and $\pm 2$ diffraction orders produced Laguerre–Gaussian modes with orbital angular momenta $\ell=\pm 2$ and $\ell=\pm 4$, respectively. Power measurements using an Ophir 3A-FS detector confirmed that the majority of the transmitted energy was contained in the zeroth and first diffraction orders, while higher orders much less visable.
The corresponding Hermite–Gaussian modes obtained after cylindrical-lens conversion are presented in Figure 13(b). The observed modal structures agree with the expected relation [6, 29]
@@
demonstrating that the grating reliably generates the targeted OAM states. Higher diffraction orders (n=3n=-3 and n=4n=-4), shown in Figure 13(c) and (d), also exhibit well-defined modal patterns, confirming correct operation of the fabricated structure across multiple orders.
[…]
Figure 13: Twisted light generated by the reflective fork grating. (a) Laguerre–Gaussian modes with the OAM $\ell=\pm 2$ and $\ell=\pm 4$ observed in the $\pm 1$ and $\pm 2$ diffraction orders. (b) Corresponding Hermite–Gaussian modes obtained after cylindrical-lens mode conversion. (c) Higher-order Laguerre–Gaussian modes with $\ell=-2,\,-4,\,-6,\,-8$ observed in the n=1,2,3,4n=-1,-2,-3,-4 diffraction orders. (d) Corresponding Hermite–Gaussian modes obtained after cylindrical-lens mode conversion.
A potential route for improving the performance of the diffractive element is to increase the spatial resolution of the grating lines. A finer grating would enable the generation of higher-quality vortex modes but would also require tighter focusing of the incident beam onto the grating surface. Overall, the experimentally measured spatial profiles in Figure 13(a) and (b) show good qualitative agreement with the simulations presented in Figure 6, confirming correct operation of the fabricated fork grating.
@@ 4.3 Experiments with SPP
The next experiment was performed using the fabricated SPP, and the resulting beam profiles are shown in Figure 14(a) and (b). The output twisted light (see Figure 14(a)) exhibits a well-defined annular structure with a pronounced central singularity. A portion of the beam energy is redistributed into several weak concentric rings surrounding the primary vortex mode, consistent with the expected behavior of high-charge SPP-generated beams.
Power measurements performed with the Ophir 3A-FS detector showed an intensity reduction of approximately $20\%$. Although this differs from the theoretical conversion efficiency of $\sim 90\%$, the obtained value is still reasonable given the surface-profile discretization, potential scattering losses, and alignment constraints in the UV. The measured twisted light structure and transverse size closely match the simulated pattern shown in Figure 8.
The Hermite–Gaussian mode (see Figure 14(b)) exhibits a slightly larger tilt compared with the simulation. This effect is attributed to a small intentional misalignment of the beam relative to the focal plane of the cylindrical lens, introduced to improve the separation and visibility of the experimentally observed mode stripes.
4.4 Experiments with binary axicons
The final series of experiments was performed using the fabricated binary axicons. The resulting quasi–Bessel beams are shown in Figure 14(c) and (d). For the axicon with the topological charge $m=3$ (see Figure 14(c)), three intensity lobes are clearly visible, consistent with the expected topological charge. At short propagation distances, these lobes partially overlap and form a characteristic Y-shaped pattern, which is typical for binary axicons due to the limited spatial separation between the initial diffraction maxima.
[…]
Figure 14: Twisted light generated by structured optical elements. (a) Laguerre–Gaussian vortex carrying the OAM $\ell=64$ generated by the SPP. (b) Corresponding Hermite–Gaussian mode of the order $N=65$ obtained after cylindrical-lens mode conversion. (c) Quasi–Bessel beam produced by the binary axicon with the topological charge $m=3$, exhibiting three intensity lobes. (d) Quasi–Bessel beam produced by the binary axicon with the topological charge $m=10$, exhibiting ten resolved lobes.
In contrast, the axicon with $m=10$ produces a well-resolved multi-lobe ring (see Figure 14(d)), in good agreement with the simulated patterns in Figure 8. Both axicons exhibit weak spiral-like background features associated with the binary phase discretization.
@@ 4.4 Experiments with binary axicons
Finally, we note that Bessel and quasi–Bessel beams cannot be converted into Hermite–Gaussian modes using a cylindrical-lens scheme, since their propagation-invariant structure does not undergo the conventional LG-to-HG transformation. For such beams, the orbital angular momentum projection simply coincides with the topological charge $m$. As shown in Ref. [26], $m$ can be obtained simply by counting the number of intensity maxima in the ring, providing a straightforward and reliable method for identifying the OAM content of quasi–Bessel beams. In practice, quasi–Bessel beams can exhibit a finite OAM bandwidth due to small contributions the nearby OAM states, but the mean topological charge still equals to $m$, as discussed in Section 3.
@@ 5 Discussion and conclusion
Figure 15: Radial profiles of the generated twisted light beams: (a)(c) – fork grating producing the OAM $\ell=2$, $\ell=4$ and $\ell=8$ modes; (d) – SPP generating a high-order LG beam with $\ell=64$; (e) – binary axicon producing a quasi–Bessel beam with the topological charge $m=10$.
We have demonstrated the generation and numerical propagation of optical beams carrying orbital angular momentum in the deep-ultraviolet range using fabricated diffractive optical elements, including fork gratings, spiral phase plates, and binary axicons. A consistent angular-spectrum–based modeling approach was used to compare the resulting transverse intensity distributions and modal properties.
Fork gratings generate Laguerre–Gaussian modes with systematically increasing ring radii and well-defined annular profiles, in agreement with the experimental radial distributions shown in Figure 15(a)–(c) and the simulations in Figure 6. Spiral phase plates provide the highest modal purity and conversion efficiency, enabling stable generation of high-order OAM states up to $\ell=64$, as illustrated in Figure 15(d) and Figure 8. Binary axicons produce low-divergence quasi–Bessel beams with segmented ring-shaped intensity profiles and a finite OAM bandwidth corresponding to a superposition of multiple OAM states, consistent with the experimental and numerical results in Figures 15(e) and 8. A comparative overview of the three diffractive elements is summarized in Table 1.
This enables the generation of high-order OAM beams up to $\ell=64$ with well-defined transverse structure, suitable for OAM transfer to relativistic electrons [36] and diagnostics in RF photoinjectors. Given that the divergence of Laguerre–Gaussian modes scales as $\sqrt{\ell}$ [24], beamlines for high-OAM injection must account for the corresponding increase of the Twiss $\beta$-function and transverse emittance [49, 11, 31]. For metallic photocathodes, spiral phase plates and axicons are particularly attractive due to their higher efficiency, modal purity, and power handling. The high modal purity of the SPP beam also facilitates OAM diagnostics based on LG–HG conversion using astigmatic optics [41], while diffraction on a triangular aperture provides a viable alternative for quasi–Bessel beams generated by axicons [32]. Overall, the presented results establish a direct connection between diffractive phase design, numerical propagation, and structured light generation in the DUV regime, relevant for photocathode-based RF photoinjectors and accelerator-driven applications.
@@
Suggested Photocathode Type Semiconductor Any Type Any Type
Table 1: Comparison of parameters for the fork grating, the spiral phase plate, and the binary axicon.
Acknowledgements
We deeply acknowledge A.A. Bogdanov for the help that made the fabrication of the spiral phase plate possible. We sincerely thank K.V. Cherepanov for invaluable assistance and advice. We are also grateful to S.S. Baturin for helpful discussions and constructive feedback. We additionally thank N.E. Sheremet for assistance with developing the numerical code used to generate the fork-grating phase masks. The axicons were simulated and fabricated at the shared research facility of Siberian Center for Synchrotron and Terahertz Radiation at the basis of the Novosibirsk Free Electron Laser, Budker Institute of Nuclear Physics SB RAS. The authors sincerely thank Maria Zhuravleva for her valuable assistance in preparing the graphical materials for this work.
This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [37].